# Hubble tension tomography: BAO vs SnIa distance tension

### v1

@@ I Introduction
    to a lower value induced by Early Dark Energy Poulin et al. (2019); Kamionkowski and Riess (2023); Simon et al. (2023); Braglia et al. (2020); Niedermann and Sloth (2020); Smith et al. (2021) (EDE), modified gravityBraglia et al. (2021); Brax et al. (2014); Adi and Kovetz (2021); Clifton et al. (2012); Lin et al. (2019); Di Valentino et al. (2016); Akrami et al. (2021) or radiationSeto and Toda (2021); Sakstein and Trodden (2020). Thus, these models exploit the degeneracy of $r_{s}$ with $H_{0}$ in the context of the observable angular scale of the CMB sound horizon $r_{s}$
@@
    The main problem of this class is that despite fine tuning of theoretical models that support them, they are only able to decrease the statistical significance of the Hubble tension and they can not fully eliminate it. In addition, they favor a higher value of matter density parameter $\Omega_{0m}$ thus worsening the $S_{8}$ tension Alestas and Perivolaropoulos (2021); Jedamzik et al. (2021); Vagnozzi (2023).
  • Late time models ($H(z)$ deformation): These models assume that there is deformation of the Hubble expansion history $H(z)$ with respect to the Planck18/$\Lambda$CDM prediction at late cosmological times ($z\lesssim 2$). This deformation shifts the value of $H_{0}$ measured by the sound horizon scale standard ruler to a value consistent with the distance ladder measurement by effectively changing the denominator of eq. (3). However, as discussed in the present analysis even this model has difficulty to fit simultaneously BAO and SnIa data.
    The main problem of this class of models is that $E(z)\equiv H(z)/H_{0}$ deformations are highly constrained by BAO and SnIa data in redshifts larger than $z\simeq 0.1$ Alestas et al. (2022); Alestas and Perivolaropoulos (2021); Alestas et al. (2020); Brieden et al. (2022); Keeley and Shafieloo (2023); Clark et al. (2021); Chen et al. (2021); Anchordoqui et al. (2022); Mau et al. (2022); Cai et al. (2022); Heisenberg et al. (2022); Vagnozzi et al. (2022); Davari and Khosravi (2022). Thus, this class of models have been put in disfavor during the last few years. Exception constitutes the $\Lambda_{\rm s}$ model which shifts the deformation to high redhifts ($z\simeq 2$) and makes it abrupt so that it has a minimal effect on late data while maintaining consistency with the distance to the CMB sound horizon at last scattering Akarsu et al. (2023a).
@@
Recent studies have investigated the redshift dependence of the Hubble tension using various probes and methods Dainotti et al. (2021); Colgáin et al. (2022); Krishnan et al. (2021). These analyses consistently find a reduction in the tension at higher redshifts, hinting at a possible low-redshift origin for the discrepancy. However, the specific redshift range and significance of this effect varies between studies, motivating further investigation.
@@ IV Conclusion and Discussion
We have explored the level of consistency between the distance moduli of an extensive up to date compilation of BAO and the Pantheon+ SnIa data as a function of redshift. Using a direct comparison of the measured luminosity distance moduli in redshift bins as well as a comparison of the $\Lambda$CDM best fit parameter values in low and high $z$ redshift bins we demonstrated the level of tension between the BAO and SnIa at various redshift bins. This discrepancy is mainly due to the well known inconsistecy of the calibrators used in each sample (recombination sound horizon vs distance ladder calibration of SnIa with Cepheids) known as the Hubble tension. Since we have used data at intermediate redshifts we were able to identify the redshift dependence of this tension as expressed through BAO and SnIa data. In this context our analysis may be viewed as a redshift tomography of the Hubble tension.
@@ Acknowledgements
This article is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). This project was also supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the ”First call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789).

### v2

@@ I Introduction
    to a lower value induced by Early Dark Energy Poulin et al. (2019); Kamionkowski and Riess (2023); Simon et al. (2023); Braglia et al. (2020a); Niedermann and Sloth (2020); Smith et al. (2021); Rezazadeh et al. (2022) (EDE), modified gravityBraglia et al. (2021); Brax et al. (2014); Adi and Kovetz (2021); Clifton et al. (2012); Lin et al. (2019); Di Valentino et al. (2016); Akrami et al. (2021); Rossi et al. (2019); Braglia et al. (2020b); Franco Abellán et al. (2023) or dark radiationSeto and Toda (2021); Sakstein and Trodden (2020); Vagnozzi (2020). Thus, these models exploit the degeneracy of $r_{s}$ with $H_{0}$ in the context of the observable angular scale of the CMB sound horizon $r_{s}$
@@
    The main problem of this class is that despite fine tuning of theoretical models that support them, they are only able to decrease the statistical significance of the Hubble tension and they can not fully eliminate it. In addition, they favor a higher value of matter density parameter $\Omega_{0m}$ thus worsening the $S_{8}$ tension Alestas and Perivolaropoulos (2021); Jedamzik et al. (2021); Vagnozzi (2023, 2021).
  • Late time models ($H(z)$ deformation): These models assume that there is deformation of the Hubble expansion history $H(z)$ with respect to the Planck18/$\Lambda$CDM prediction at late cosmological times ($z\lesssim 2$) Li and Shafieloo (2019); Pan et al. (2020); Panpanich et al. (2021); Di Valentino et al. (2020a, b); Li and Shafieloo (2020); Clark et al. (2021); Di Valentino et al. (2020c). This deformation shifts the value of $H_{0}$ measured by the sound horizon scale standard ruler to a value consistent with the distance ladder measurement by effectively changing the denominator of eq. (3). However, as discussed in the present analysis even this model has difficulty to fit simultaneously BAO and SnIa data.
    The main problem of this class of models is that $E(z)\equiv H(z)/H_{0}$ deformations are highly constrained by BAO and SnIa data in redshifts larger than $z\simeq 0.1$ Alestas et al. (2022); Alestas and Perivolaropoulos (2021); Alestas et al. (2020); Brieden et al. (2022); Keeley and Shafieloo (2023); Clark et al. (2021); Chen et al. (2021); Anchordoqui et al. (2022); Mau et al. (2022); Cai et al. (2022); Heisenberg et al. (2022); Vagnozzi et al. (2022); Davari and Khosravi (2022); Di Valentino et al. (2020b); Gómez-Valent et al. (2024). Thus, this class of models have been put in disfavor during the last few years. Exception constitutes the $\Lambda_{\rm s}$ model which shifts the deformation to high redhifts ($z\simeq 2$) and makes it abrupt so that it has a minimal effect on late data while maintaining consistency with the distance to the CMB sound horizon at last scattering Akarsu et al. (2023a).
@@
Recent studies have investigated the redshift dependence of the Hubble tension using various probes and methods Dainotti et al. (2021); Colgáin et al. (2022); Krishnan et al. (2021a); Jia et al. (2023); Krishnan et al. (2020, 2021b). These analyses consistently find a reduction in the tension at higher redshifts, hinting at a possible low-redshift origin for the discrepancy. However, the specific redshift range and significance of this effect varies between studies, motivating further investigation.
@@ II.2 Distance moduli from BAO
The data in Table 1 are from anisotropic BAO analyses, they have been derived using a fiducial model to convert the observed data into physical distances, and then isolating the angular component ($D_{M}$) to compute $D_{M}/r_{d}$. The distinction between anisotropic BAO analyses and 2D BAO data may be described as follows:
  • Anisotropic BAO analyses incorporate the full three-dimensional distribution of galaxies. These analyses take into account both the radial and transverse components simultaneously, often using a fiducial cosmological model to convert observed angles and redshifts into physical distances. This allows for a more detailed understanding of the BAO signal and its anisotropies, but it introduces model dependencies in the process.
  • In 2D BAO analyses, the BAO signal is examined in a two-dimensional plane, often focusing separately on the transverse (angular) and radial (line-of-sight) directions. This method typically avoids using a fiducial cosmological model to convert angular separations and redshifts directly into physical distances. Instead, it keeps the analysis in observational units (angles and redshifts), which can then be used to infer cosmological parameters in a somewhat model-independent way. This approach is designed to minimize assumptions about the underlying cosmology when extracting the BAO signal.
@@ IV Conclusion and Discussion
We have explored the level of consistency between the distance moduli of an extensive up to date compilation of BAO and the Pantheon+ SnIa data as a function of redshift. Using a direct comparison of the measured luminosity distance moduli in redshift bins as well as a comparison of the $\Lambda$CDM best fit parameter values in low and high $z$ redshift bins we demonstrated the level of tension between the BAO and SnIa at various redshift bins. This discrepancy is mainly due to the well known inconsistency of the calibrators used in each sample (recombination sound horizon vs distance ladder calibration of SnIa with Cepheids) known as the Hubble tension. Since we have used data at intermediate redshifts we were able to identify the redshift dependence of this tension as expressed through BAO and SnIa data. In this context our analysis may be viewed as a redshift tomography of the Hubble tension.
@@
While there may be correlation among some of the BAO points of our sample, this covariance is in most cases not available in the literature. Thus, by including most available BAO data, we chose to maximize information rather than independence. This approximation may have introduced some bias in our results. Since our sample is not very extensive we estimate that this bias is subdominant and does not change the main features of our results.
@@ Acknowledgements
We thank Adrià Gómez-Valent, Sunny Vagnozzi, Santi Avila and Eoin Colgain for useful comments. This article is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). This project was also supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the ”First call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789).